Problem 15-4A (Part Level Submission) Agassi Company uses a job order cost system in each of its three manufacturing departments. Manufacturing overhead is applied to jobs on the basis of direct labor cost in Department D, direct labor hours in Department E, and machine hours in Department K. In establishing the predetermined overhead rates for 2017, the following estimates were made for the year. Department D E K Manufacturing overhead $854,000 $1,500,000 $1,020,000 Direct labor costs $1,220,000 $1,875,000 $675,000 Direct labor hours 150,000 125,000 60,000 Machine hours 600,000 750,000 120,000 During January, the job cost sheets showed the following costs and production data. Department D E K Direct materials used $210,000 $189,000 $117,000 Direct labor costs $180,000 $165,000 $56,250 Manufacturing overhead incurred $148,500 $186,000 $118,500 Direct labor hours 12,000 16,500 5,250 Machine hours 51,000 67,500 10,390 (a) Your answer is correct. Compute the predetermined overhead rate for each department. (Round answers to 2 decimal places, e.g. 10.25.) Overhead rate Department D % Department E $ per direct labor hour Department K $ per machine hour Click if you would like to Show Work for this question: Open Show Work Show Solution Show Answer Link to Text Attempts: 1 of 3 used (b) Compute the total manufacturing costs assigned to jobs in January in each department. (Round answers to 0 decimal places, e.g. 2,500.) Manufacturing Costs Department D $ Department E $ Department K $
Answer #1
PARTICULAR DEPARTMENT D DEPARTMENT E DEPARTMENT K
PREDETERMINED
1. MANUFACTURING OVERHEAD $ 854000 $ 1500000 $ 1020000
2. DIRECT LABOR COST $ 1220000 $ 1875000 $ 675000
3. DIRECT LABOR HOUR 150000 125000 60000
4. MACHINE HOUR 600000 750000 120000
ACTUAL FOR JAN.
5. DIRECT MATERIAL USED $ 210000 $ 189000 $ 117000
6. MANUFACTURING OVERHEAD $ 148500 $ 186000 $ 118500
7. DIRECT LABOR COST $ 180000 $ 165000 $ 56250
8. DIRECT LABOR HOUR 12000 16500 5250
9. MACHINE HOUR 51000 67500 10390
(a) PREDETERMINED OVERHEAD RATE
BASIS DIRECT LABOR COST DIRECT LABOR HOUR MACHINE HOUR
= 1 /2 x 100 = 1 / 3 = 1 / 4
= 854000 / 1220000 x 100 = 70 % = 1500000 / 125000 = $ 12.00 PER DIRECT LABOR HOUR = 1020000 / 120000 = $ 8.50 PER MACHINE HOUR
(b) TOTAL MANUFACTURING COST ASSIGNED TO JOBS IN JAN. IN EACH DEPARTMENT = 7 x (a) = 8 x (a) = 9 x (a)
= 180000 x 70% = $126000 = 16500 x 12.00 = $198000 = 10390 x 8.5 = $88315